### v3

# Estimating network-mediated causal effects via principal components network regression

Abstract
We develop a method to decompose causal effects on a social network into an indirect effect mediated by the network, and a direct effect independent of the social network. To handle the complexity of network structures, we assume that latent social groups act as causal mediators. We develop principal components network regression models to differentiate the social effect from the non-social effect. Fitting the regression models is as simple as principal components analysis followed by ordinary least squares estimation. We prove asymptotic theory for regression coefficients from this procedure and show that it is widely applicable, allowing for a variety of distributions on the regression errors and network edges. We carefully characterize the counterfactual assumptions necessary to use the regression models for causal inference, and show that current approaches to causal network regression may result in over-control bias. The structure of our method is very general, so that it is applicable to many types of structured data beyond social networks, such as text, areal data, psychometrics, images and omics.
Keywords: causal mediation, latent mediators, network regression, principal components regression, random dot product graph, spectral embedding
1 Introduction
Recent years have seen a concerted effort to study causal effects on networks, motivated by striking claims about contagions in social networks (Christakis and Fowler, 2007). One of the key ideas to emerge from this push is the need to account for clustering in networks (Shalizi and Thomas, 2011). Sociologists have long known that people in social networks are mostly connected to other people like themselves, which is often expressed informally as “birds of a feather flock together,” and more formally called “homophily”. To identify and estimate causal effects in social settings, it is thus fundamental to model how social groups form in networks, as well any downstream effects of social group membership. This is challenging, as social groups in a network are typically unobserved.
To account for unobserved social structure in networks, Shalizi and Thomas (2011) proposed using latent space models, where each node has an embedding that determines its propensity to connect with other nodes. More recently, McFowland and Shalizi (2021) showed that certain types of causal effects on networks can be estimated by controlling for latent node embeddings in linear network regression models. There is now a rapidly growing literature investigating how embeddings can be used for causal inference on networks (Paul et al., 2022a; Veitch et al., 2019, 2020; Louizos et al., 2017; Guo et al., 2020; Chu et al., 2021; Guo et al., 2020; Cristali and Veitch, 2022; Chen et al., 2022; Ogburn et al., 2022; Liu and Tchetgen Tchetgen, 2022; O’Malley et al., 2014; Leung, 2019; Egami and Tchetgen Tchetgen, 2021; Ogburn et al., 2022, 2020).
We make two contributions to this literature. First, we develop broad semi-parametric theory for network regression, showing how to incorporate node embeddings into linear regression. These network regression models are useful for causal inference on networks, as they allow practitioners to account for homophily, but they are also of substantial independent interest as an observation model. Similar models have previously appeared in Li et al. (2019); Le and Li (2022); Fosdick and Hoff (2015); He and Hoff (2019) and in concurrent work Nath et al. (2023); Chang and Paul (2024). Compared to these approaches, we require substantially weaker assumptions on the edge distribution in the network and the error distribution in the regression models. We allow weighted (i.e., real-valued) edges and allow edges to be observed with noise. Edges in the network and regression errors can be sampled from any distribution that satisfies a sub-gamma tail bound. Further, our regression models allow for heteroscedastic errors. Altogether, our results show estimating network embeddings with principal components analysis, together with ordinary least squares to estimate regression coefficients, is applicable under broad semi-parametric conditions. This contrasts with previous approaches that have largely focused on specific parametric models.
Our second contribution is to characterize how latent node-level variables (i.e., embeddings) can act as causal mediators (Imai et al., 2010). To date, causal methods for network data have mainly focused on homophily as a confounding factor. In contrast, we articulate the causal mechanisms involved when treatments influence latent social group membership. This is perhaps best illustrated by an example. Suppose we are interested in understanding and reducing adolescent smoking (Di Maria et al., 2022; Michell and Amos, 1997). Understanding why adolescents smoke is challenging, as smoking is both a sexually differentiated behavior and a social behavior. To study the effect of sex on adolescent smoking, we decompose the effect of sex into a direct effect independent of the network and an indirect effect that operates through the network. For instance, sex may directly cause higher cigarette consumption through sexed expectations about smoking. Sex may also have an indirect effect mediated by the network: adolescents may prefer to form sex-homophilous friendships, and social norms about the acceptability of smoking may vary across friend groups. That is, sex may influence an adolescent’s social circumstances, which in turns causes a change in smoking behavior. These effects correspond to natural direct and indirect effects, specialized to the network setting by treating latent community memberships as mediators.
To estimate network-mediated effects, we propose an approach based on a generalization of the random dot product graph (Athreya et al., 2018; Levin et al., 2022) and principal components network regression (Le and Li, 2022; Cai et al., 2021; Paul et al., 2022a; McFowland and Shalizi, 2021; Upton and Carvalho, 2017). To represent the mediating effect of the network, we embed the network into a low-dimensional Euclidean space via a singular value decomposition (Sussman et al., 2014). We then use the embeddings to develop two network regression models: (1) an outcome model that characterizes how nodal outcomes vary with nodal treatment, controls, and position in latent space; and (2) a mediator model that characterizes how latent positions vary with nodal treatment and controls. We estimate the regressions using ordinary least squares and Huber-White robust standard errors, and prove that the resulting coefficients are asymptotically normal under general semi-parametric conditions. Once estimation is complete, the coefficients from the outcome model can be used to estimate the direct effect of treatment, and the coefficients from the outcome model and the mediator model can be combined to estimate the indirect effect of treatment (VanderWeele and Vansteelandt, 2014). These estimators are essentially the well-known product-of-coefficients estimators (VanderWeele and Vansteelandt, 2014; Nguyen et al., 2021), but using the network embedding as a mediator. We anticipate that our causal estimators will be accessible to many social and natural scientists, as they are based on familiar product-of-coefficients mediation techniques. Importantly, the mediation estimands we consider in this work are distinct from peer effects such as contagion and interference, explored elsewhere in the literature (Ogburn et al., 2020, 2022). Our mediational estimands measure the effect of meso-level social structures, in contrast to more micro-level peer influence. In Section 2.5 we explore the relationship between these mechanisms.
While the product-of-coefficient estimator is familiar to many, it relies on strong functional form assumptions. We want to emphasize that our characterization of latent mediation is valuable even to readers who suspect these assumptions are too strong to hold in practice. Following McFowland and Shalizi (2021), practitioners and theorists are increasingly using linear regressions for causal inference on networks, and careful charactizations of the causal structure of these models is increasingly important (Paul et al., 2022a; McFowland and Shalizi, 2021; Veitch et al., 2019, 2020; Louizos et al., 2017; Guo et al., 2020; Chu et al., 2021; Guo et al., 2020; Cristali and Veitch, 2022; Chen et al., 2022; Ogburn et al., 2022; Liu and Tchetgen Tchetgen, 2022; O’Malley et al., 2014; Leung, 2019; Egami and Tchetgen Tchetgen, 2021; Ogburn et al., 2022, 2020). In our development of the product-of-coefficients estimator, we correct the misconception that homophily can only have a confounding effect, we describe the over-control bias that can be induced by causal misspecification, and we document a network dataset where causal misspecification leads to substantially misleading results. Indeed, this work was originally motivated by a misinterpretation of coefficients in a principal components network regression model.
We expect our estimators to have significant causal and non-causal applications to outside of the network setting. Similar methods for low-rank data have been employed, for example, on spatial networks (Gilbert et al., 2021; Tiefelsdorf and Griffith, 2007; Doreian, 1981; Ord, 1975), text data (Keith et al., 2021; Veitch et al., 2020; Gerlach et al., 2018), psychometric surveys (Freier et al., 2022; Thurstone, 1947), imaging data (Zhao et al., 2020; Levin et al., 2022), and omics panels (Listgarten et al., 2010; Alter et al., 2000). Indeed, we apply our method to psychometric data, in addition to our running social network example. In data applications, we find that adolescent girls end up smoking more than adolescent boys primarily due to an indirect network effect. This suggests that public health interventions might want to disrupt or alter social group formation. In an application to psychometrics data, we find that a meditation app reduces anxiety primarily by helping study participants defuse (i.e., step back and examine) from their emotions and by helping them feel less lonely. Our findings suggest that the meditation program might improve mental health outcomes by replacing meditation modules aimed at increasing a sense of meaning in life with additional modules targeting defusion or loneliness.
Our work is related to several extant lines of research. Past authors have heuristically proposed regression estimators for latent mediation in networks under the Hoff model (Hoff et al., 2002), albeit without asymptotic theory or much elaboration of causal mechanisms (Di Maria et al., 2022; Liu et al., 2021; Che et al., 2021). These ideas are related to work on non-parametric mediation analysis (Tchetgen Tchetgen and Shpitser, 2012; Farbmacher et al., 2022; Fulcher et al., 2020; Zheng and van der Laan, 2012) and semi-parametric methods for hidden mediators such as Cheng et al. (2022), as well as methods for proximal mediation such as Dukes et al. (2021) and Ghassami et al. (2021). Similarly, Sweet (2019); Sweet and Adhikari (2022); Guha and Rodriguez (2021); Zhao et al. (2022) consider mediation in networks, albeit treating entire networks as mediators, an approach that requires observing an entire network per unit of analysis.
Beyond causal considerations, this paper proposes semi-parametric methods for principal components network regression. Similar results have previously been established in more restrictive parametric settings. Le and Li (2022) and Paul et al. (2022a) show a related result in binary networks, under the assumption that regression errors are Gaussian, and Cai et al. (2021) considers Gaussian networks with a one-dimensional latent space. We substantially relax these assumptions to allow for latent spaces of arbitrary dimension and to permit far more general edge and regression error distributions. The regression models that we propose are related to several other forms of regression studied within the network literature, such as classic spatial and econometric interference models (Land and Deane, 1992; Manski, 1993) and regressions with network-coherence penalties (Li et al., 2019). For a review, we refer the interested reader to Le and Li (2022). Also related is a large body of work on network association testing, such as Ehrhardt and Wolfe (2019); Fredrickson and Chen (2019); Lee et al. (2019); Gao et al. (2022); Su et al. (2020).
Notation
For a matrix $A$, let $\left\lVert A\right\rVert,\left\lVert A\right\rVert_{F}$ and $\left\lVert A\right\rVert_{2,\infty}$ denote the spectral, Frobenius, and two-to-infinity norms, respectively. We write $A^{\dagger}$ for the Moore-Penrose pseudoinverse, $A_{i\cdot}$ for the $i$-th row, $A_{\cdot j}$ for the $j$-th column, and $\operatorname*{vec}(A)$ for the column-wise vectorization of $A$, i.e., $\operatorname*{vec}(A)=(A_{\cdot 1}^{T},A_{\cdot 2}^{T},\dots,A_{\cdot n}^{T})^{T}$ for a matrix with $n$ columns. We use $\otimes$ to denote the Kronecker product. We write $[n]$ to denote the set $\{1,2,\dots,n\}$. $\mathbb{O}_{d}$ denotes the set of $d\times d$ orthogonal matrices. When we define a new symbol inline, we use $\equiv$. We use standard Landau notation, e.g., $\mathcal{O}\left(a_{n}\right)$ and $o\left(a_{n}\right)$ to denote growth rates, as well as the probabilistic variants $\mathcal{O}_{p}\left(a_{n}\right)$ and $o_{p}\left(a_{n}\right)$. For example, $g(n)=\mathcal{O}\left(f(n)\right)$ means that for some constant $C>0$, $|g(n)|<Cf(n)$ for all suitably large $n$. In proofs, $C$ denotes a constant not depending on the number of vertices $n$, whose precise value may change from line to line, and occasionally within the same line.
2 The causal structure of mediation in latent spaces
A central goal of this paper is to formalize and estimate natural direct and indirect effects as mediated by latent positions in a network. Briefly, the idea is that node-level treatments can effect formation of social groups, and that membership in social groups can further influence nodal outcomes.
2.1 Motivating example: social and non-social elements of teenage smoking
Let us return to the adolescent smoking example mentioned previously, which is based on the Teenage Friends and Lifestyle Study, reported in Michell and West (1996), Michell and Amos (1997), Michell (1997), and Michell (2000b). The Teenage Friends and Lifestyle Study collected three waves of survey data in a secondary school in Glasgow, beginning in January 1995. Students in the study filled out a questionnaire about their lifestyle and risk-taking behaviors, including alcohol, tobacco and drug use, and additionally were asked to list six of their friends. Beyond this quantitative data, researchers also conducted in-depth qualitative work, investigating the social dynamics of the school through focus groups, classroom observations, and interviews.
In the Glasgow data, as in prior investigations, researchers found that adolescent girls smoked more than adolescent boys. Michell and West (1996) and Michell (2000b) proposed two distinct effects on smoking: one about athletic expectations, and the other about social influence. Michell (2000b) proposed that athletic expectations affected smoking as follows: both boys and girls were subject to general societal pressure to smoke, but adolescent boys were subject to additional pressure to be good at sports (Michell and West, 1996). Since smoking was generally accepted as reducing athletic performance, this influenced adolescent boys to abstain from smoking in favor of pursuing athletic social capital (Michell, 2000b).
The second proposed effect was a social effect. Smoking is known to be a collective behavior, where group members typically all partake or all abstain (Michell, 2000b). Researchers found evidence of this form of group decision-making for various risk-taking behaviors in the Glasgow study, including tobacco consumption. This group-level decision-making led to differentiated behavior amongst adolescent boys and girls because adolescent friendships are highly sex-homophilous until the onset of puberty (Mehta and Strough, 2009). Taken all together, this meant that the effect of sex on smoking was mediated by friend group membership: adolescent friend group membership was heavily determined by sex, and smoking was heavily determined by friend group. More precisely, Michell (2000b) found that smoking was mostly concentrated in friend groups composed of popular girls, unpopular students, and trouble-makers: “risk taking behaviour was heavily polarized within social categories so that, for instance, groups of individuals (and their peripherals) were in general either risk-taking or non-risk-taking […] generally groups (and their peripherals) were either all boys or all girls.”
Figure 1: Directed friendships in a secondary school in Glasgow, reported in the Teenage Friends and Lifestyle Study (wave 1). Each node represents one student. An arrow from node $i$ to node $j$ indicates student $i$ claimed student $j$ as a friend. Node size is proportional to in-degree.
The goal of this paper is to formalize a causal model for network-linked data that allows us to quantify and estimate causal effects like those proposed in the example above, a setting that we refer to as homophilous mediation. The first mechanism corresponds to a direct effect, independent of social considerations, and the second mechanism corresponds to an indirect effect, with friend group membership mediating the causal effect of sex on smoking. In the Glasgow study, it seems intuitive that sex exerted both a direct and indirect effect on tobacco consumption, driving adolescent girls to smoke more than adolescent boys. Our formalization, developed in Sections 2.2 and 2.3, also accommodates observed confounders influencing both the direct and the indirect pathways. For example, some students in the Glasgow study reported attending church, which has the potential to influence both their social group and attitudes about tobacco.
2.2 Causal mediation in latent social spaces
To formalize a model for mediated effects on a network, we begin by introducing some notation. We assume that we have a network with $n$ nodes (corresponding to our experimental units), which we label according to the integers $[n]=\{1,2,\dots,n\}$. Let $T\in\{0,1\}^{n}$ be a vector of observed binary treatment indicators for nodes $i\in[n]$, and let $Y\in\mathbb{R}^{n}$ be a vector of observed node-level outcomes. Let $X\in\mathbb{R}^{n\times d}$ be a matrix describing the unobserved “friend groups” of each node (to be formalized below), with the row vector $X_{i\cdot}\in\mathbb{R}^{1\times d}$ encoding the friend group of node $i$ for each $i\in[n]$. Similarly, let $C\in\mathbb{R}^{n\times p}$ be a matrix of observed confounders, with the row vector $C_{i\cdot}\in\mathbb{R}^{1\times p}$ denoting the confounders associated with node $i$. Lastly, let $A_{ij}\in\mathbb{R}$ denote the strength of the friendship between node $i$ and node $j$. For simplicity, our model takes friendships to be symmetric, such that $A_{ij}=A_{ji}$ for all $i,j\in[n]$, but this condition can be relaxed to allow for directed networks (see Section 5).
To define causal estimands, we introduce notation for the necessary counterfactual quantities, defined in the sense of structural causal models (Pearl, 2009).
Definition 1
Let $Y_{i}(t)$ be the counterfactual value of the outcome measured for the $i^{th}$ node when $T_{i}$ is set to $t$. Similarly, let $Y_{i}(t,x)$ be the counterfactual value of the $i^{th}$ outcome when $T_{i}$ is set to $t$ and $X_{i\cdot}$ is set to $x$, and let $X_{i\cdot}(t)$ be the counterfactual value of the mediator $X_{i\cdot}$ when $T_{i}$ is set to $t$. The average treatment effect, natural direct effect and natural indirect effect are defined, respectively, as
$\Psi_{\mathrm{ate}}\left(t,t^{*}\right)$ $=\mathbb{E}\left[Y_{i}(t)-Y_{i}(t^{*})\right],$
$\Psi_{\mathrm{nde}}\left(t,t^{*}\right)$ $=\mathbb{E}\left[Y_{i}(t,X_{i\cdot}(t^{*}))-Y_{i}(t^{*},X_{i\cdot}(t^{*}))\right],\text{ and }$
$\Psi_{\mathrm{nie}}\left(t,t^{*}\right)$ $=\mathbb{E}\left[Y_{i}(t,X_{i\cdot}(t))-Y_{i}(t,X_{i\cdot}(t^{*}))\right].$
Note that the average treatment effect decomposes into the sum of the natural direct effect and the natural indirect effect: $\Psi_{\mathrm{ate}}\left(t,t^{*}\right)=\Psi_{\mathrm{nde}}\left(t,t^{*}\right)+\Psi_{\mathrm{nie}}\left(t,t^{*}\right)$. We interpret all three estimands following Chapter 2 of VanderWeele (2015): the average treatment effect $\Psi_{\mathrm{ate}}$ describes how much the outcome $Y_{i}$ would change on average over all nodes $i\in[n]$ if the treatment $T_{i}$ were changed from $T_{i}=t$ to $T_{i}=t^{*}$. The natural direct effect describes how much the outcome $Y_{i}$ would change on average over all nodes $i\in[n]$ if the exposure $T_{i}$ were set at level $T_{i}=t^{*}$ versus $T_{i}=t$ but for each individual the mediator $X_{i\cdot}$ were kept at the level it would have taken for that individual, had $T_{i}$ been set to $t^{*}$. The natural indirect effect describes how much the outcome $Y_{i}$ would change on average over all nodes $i=[n]$ if the exposure were fixed at level $T_{i}=t^{*}$ but the mediator $X_{i\cdot}$ were changed from the level it would take under $T_{i}=t$ to the level it would take under $T_{i}=t^{*}$.
In slightly more plain language, we can interpret the natural direct effect as capturing the effect of the exposure on the outcome when the mediating pathway is disabled. In the smoking example discussed in Section 2.1, this would correspond to the effect of sexed expectations alone. Similarly, we can interpret the natural indirect effect as capturing the effect of the exposure on the outcome that operates by changing the mediator while keeping treatment fixed (VanderWeele, 2015). In the smoking example, this corresponds to the effect of sex on friend group, and then friend group on smoking behavior. The total effect of sex on smoking is the sum of the effects from sexed athletic expectations and friend group pressures.
In order to estimate counterfactual quantities, we must make identifying assumptions to relate unobserved counterfactual quantities to the observable data. Identification is implied by the properties of consistency, sequential ignorability, and positivity (Imai et al., 2010), which are standard sufficient conditions within the causal inference literature.
Assumption 1 (Non-parametric Identification of Natural Direct and Indirect Effects)
The random variables $(Y_{i},Y_{i}(t,x),X_{i\cdot},X_{i\cdot}(t),C_{i\cdot},T_{i})$ are independent over $i\in[n]$ and obey the following three properties.
  1. 1.
    Consistency:
    if $T_{i}=t$, then $X_{i\cdot}(t)=X_{i\cdot}$ with probability 1, and
    if $T_{i}=t$ and $X_{i\cdot}=x$, then $Y_{i}(t,x)=Y_{i}$ with probability 1
  2. 2.
    Sequential ignorability:
    $\{Y_{i}(t^{*},x),X_{i\cdot}(t)\}\perp\!\!\!\perp T_{i}\mid C_{i\cdot}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \{Y_{i}(t^{*},x)\}\perp\!\!\!\perp X_{i\cdot}\mid T_{i}=t,C_{i\cdot}$
  3. 3.
    Positivity:
    $\mathbb{P}\left(x\,\middle|\,T_{i},C_{i\cdot}\right)$ $>0\text{ for each }x\in\operatorname*{supp}(X_{i\cdot})$
    $\mathbb{P}\left(t\,\middle|\,C_{i\cdot}\right)$ $>0\text{ for each }t\in\operatorname*{supp}(T_{i})$
Roughly speaking, a sufficient condition for natural direct and indirect effects to be non-parametrically identified is that the observed controls $C_{i\cdot}$ contain all confounders of the exposure-outcome ($T_{i}\to Y_{i}$), the exposure-mediator ($T_{i}\to X_{i\cdot}$) and the mediator-outcome ($X_{i\cdot}\to Y_{i}$) relationships. One structural causal model satisfying these requirements is given in Figure 2. Note that, crucially, our model differs from traditional mediation models because the mediators (i.e., the group memberships $X_{i\cdot}$) are unobserved.
Three assumptions are particularly important from a counterfactual perspective. As in tabular settings, the sequential ignorability assumption is strong and may not hold due to mediator-outcome confounding. Two other concerns are more specific to the network setting. In particular, positivity may be a problem if friend groups are highly homophilous. That is, conditional on treatment and controls, some regions of the latent space might have zero probability mass. In the context of the adolescent smoking example, this is potentially an issue, as friend groups are highly sexually homophilous. Since empirical networks often exhibit high degrees of homophily, positivity violations may present a larger challenge in network settings than in non-network settings. A third crucial assumption is that there are not peer effects such as contagion or interference, a topic we discuss in detail in Section 2.5.
Figure 2: A directed acyclic graph (DAG) representing the causal pathways of latent mediation in a network with two nodes called $i$ and $j$. Each node in the figure corresponds to a random variable, and edges indicate which random variables may cause which other random variables. We are interested in the causal effect of $T_{i}$ on $Y_{i}$, as mediated by the latent position $X_{i\cdot}$.
2.3 Semi-parametric latent space structure
The counterfactual model we have presented thus far is a standard causal model for mediation, with the exception that we have not clarified the role of $X_{i\cdot}$, which we claimed should correspond to a latent measure of social group membership. We now fix ideas about $X_{i\cdot}$ and clarify the details of the network model by developing a statistical model for low-rank matrices. This low-rank model can be thought of as a generalization of the random dot product model (Athreya et al., 2018; Nickel, 2006; Bonato and Chung, 2007) to networks with weighted (i.e., non-binary) edges.
Describing (possibly) weighted edges requires a brief technical pre-requisite.
Definition 2
Let $Z$ be a mean-zero random variable with cumulant generating function $\psi_{Z}(t)=\log\mathbb{E}\left[e^{tZ}\right]$.
  1. 1.
    $Z$ is $\nu$-sub-Gaussian for $\nu>0$ if $\psi_{Z}(t)\leq t^{2}\nu/2$ for all $t\in\mathbb{R}$.
  2. 2.
    $Z$ is $(\nu,b)$-sub-gamma for $\nu,b\geq 0$ if $\psi_{Z}(t)\leq\frac{t^{2}\nu}{2(1-bt)}$ and $\psi_{-Z}(t)\leq\frac{t^{2}\nu}{2(1-bt)}$ for all $t<1/b$.
The class of sub-gamma distributions is broad, and includes as special cases the Bernoulli, Poisson, Exponential, Gamma, and Gaussian distributions, as well as any sub-Gaussian or squared sub-Gaussian distribution, and all bounded distributions (see Boucheron et al., 2013, for a detailed treatment). Our primary assumption on the network structure is that the edges are sampled according to sub-gamma distributions and the network is low-rank in expectation (Boucheron et al., 2013; Tropp, 2015).
Assumption 2 (Sub-gamma network)
Let $A\in\mathbb{R}^{n\times n}$ be a random symmetric matrix, such as the adjacency matrix of an undirected graph. Let $P=\mathbb{E}\left[A\,\middle|\,X\right]=XX^{T}$ be the expectation of $A$ conditional on $X\in\mathbb{R}^{n\times d}$, which has independent and identically distributed rows $X_{1\cdot},\dots,X_{n\cdot}$. The matrix $P$ has $\operatorname*{rank}\left(P\right)=d$ and is positive semi-definite with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}>0=\lambda_{d+1}=\cdots=\lambda_{n}$. Conditional on $X$, the upper-triangular elements of $A-P$ are independent $(\nu_{n},b_{n})$-sub-gamma random variables.
This characterization of the network structure is very general and encompasses a range of popular network models as special cases. The most important example of the sub-gamma model for our purposes is the random dot product graph.
Example 1 (Random Dot Product Graph; Athreya et al. (2018))
Let $F$ be a distribution on $\mathbb{R}^{d}$ such that $x^{T}y\in[0,1]$ whenever $x,y\in\operatorname*{supp}F$. Draw $X_{1\cdot},X_{2\cdot},\dots,X_{n\cdot}$ i.i.d. according to $F$, and collect these $n$ points in the rows of $X\in\mathbb{R}^{n\times d}$. Conditional on $X$, the edges of graph $G$ are generated independently, with probability of an edge $i\sim j$ given by $X_{i\cdot}X_{j\cdot}^{T}$. That is, conditional on $X$, the entries of the symmetric adjacency matrix $A$ above the diagonal are independent with $A_{ij}\sim\operatorname*{Bernoulli}(X_{i\cdot}X_{j\cdot}^{T})$. Then we say that $A$ is distributed according to a random dot product graph with latent position distribution $F$ and write $(A,X)\sim\operatorname{RDPG}(F,n)$.
Under the random dot product graph, each node in a network is associated with a latent vector, and these latent vectors characterize propensities to form edges with other nodes. Specifically, nodes close to each other in latent space are more likely to form connections, and nodes far apart are unlikely to form connections. When nodes cluster in the latent space, the result is that edges in the network are more likely to form between nodes with similar latent characteristics. This manifests as homophily in the resulting network.
Another particularly important sub-gamma model is the stochastic blockmodel, which is in fact a sub-model of the random dot product graph.
Example 2 (Stochastic Blockmodel)
The stochastic blockmodel (SBM; Holland et al., 1983) is a model of community membership, in which each vertex is assigned to a community (sometimes called a “block”). Conditional on assignments to communities, edges are generated independently between every pair of vertices in the network, and the probability of forming an edge between vertices $i$ and $j$ depends only on the community memberships of nodes $i$ and $j$.
Let $B\in[0,1]^{d\times d}$ denote a fixed, positive semi-definite matrix of inter-block edge probabilities and let $Z_{i\cdot}\in\mathbb{R}^{d}$ be a vector encoding the block membership of node $i$. Conditional on the matrix $B\in[0,1]^{d\times d}$ and on the block memberships encoded by the rows of $Z\in\{0,1\}^{n\times d}$, the behavior of the stochastic blockmodel is characterized by
$\mathbb{P}\left(A_{ij}=1\,\middle|\,Z,B\right)=Z_{i\cdot}BZ_{j\cdot}^{T}.$
The basic stochastic blockmodel, as introduced in Holland et al. (1983), forces each node to belong to exactly one block. That is, it requires $Z_{i\cdot}\in\{0,1\}^{d}$ to have exactly one entry equal to one. The degree-corrected stochastic blockmodel (Karrer and Newman, 2011) relaxes this restriction by giving each node a “degree-heterogeneity” parameter that encodes a node’s propensity to form edges. This is equivalent to requiring $Z_{i\cdot}\in\mathbb{R}_{+}^{d}$, where $d-1$ entries of $Z_{i\cdot}$ are still zero. The overlapping stochastic blockmodel (Latouche et al., 2011) allows $Z_{i\cdot}\in\{0,1\}^{d}$ with no additional restrictions, such that nodes can belong to multiple blocks. The mixed membership stochastic blockmodel (Airoldi et al., 2008) restricts $Z_{i\cdot}$ to lie on the $d-1$ dimensional simplex, such that nodes can have partial membership in multiple blocks, but their total block participation must sum to one. The overlapping and mixed-membership variants may also be extended to include degree-correction, in which case $Z_{i\cdot}$ can be a nearly arbitrary vector in +$\mathbb{R}^{d}_{+}$ (Jin et al., 2024; Zhang et al., 2020).
To express the stochastic blockmodel in more general sub-gamma form, take $X=ZB^{1/2}$. Thus, the latent positions $X$ encode (1) block participation as characterized by $Z$, (2) degree-adjustment (i.e., popularity) captured by the row scales of $Z$, and (3) intra-block edge formation propensities contained in $B$.
Before moving on to describe how the latent positions $X$ are related to other network covariates, some remarks about the general sub-gamma model are warranted. First, the sub-gamma model allows for edges to be observed with noise.
Example 3 (Noisily Observed Random Dot Product Graph)
Let $\left(\mathscr{A},X\right)\sim\operatorname{RDPG}(F,n)$ and let $\{E_{ij}:1\leq i<j\leq n\}$ be independent, mean-zero sub-gamma random variables for $1\leq i\leq j\leq n$. For example, $E_{ij}$ might correspond to (centered) Gaussian or Bernoulli noise. Then the network $A_{ij}=\mathscr{A}_{ij}+E_{ij}$ satisfies Assumption 2, as the sum of sub-gamma random variables remains sub-gamma.
We note that all our results presented below can be extended to asymmetric $P$, rectangular $P$, and $P$ with negative eigenvalues. The assumption that $P$ is symmetric and positive semi-definitive is primarily to simplify notation, and our proofs can be extended to the general case using the techniques of Rubin-Delanchy et al. (2022) and Rohe and Zeng (2023). Thus, the sub-gamma model can be extended to handle bipartite and directed graphs (Qing and Wang, 2021; Rohe et al., 2016). Other natural extensions include models such as Gaussian mixtures with identity covariance, latent Dirichlet allocation (Rohe and Zeng, 2023; Blei et al., 2003), topic models (Gerlach et al., 2018), and psychometric factor models (Thurstone, 1947, 1934).
It is also important to note that, like the random dot product graph, the sub-gamma model considered here is subject to orthogonal non-identifiability. Since $P=XX^{T}=(XQ)(XQ)^{T}$ for any $d\times d$ orthogonal matrix $Q$, the latent positions $X$ are only identifiable up to an orthogonal transformation (see Athreya et al. 2018 for further discussion). Luckily, this non-identifiability of latent positions does not influence identifiability of the natural direct and indirect effects.
Lastly, under the sub-gamma model, the latent positions $X$ are sufficient for $A$. That is, the nodal covariates $T_{i},C_{i\cdot}$ and $Y_{i}$ do not directly influence the formation of edges of $A_{ij}$. The covariates $T_{i}$ and $C_{i\cdot}$ can influence edge formation, but only via the intermediary $X$. Relaxing this constraint is an interesting topic for future work, but requires different estimators for the latent positions $X$ than those we consider here (Mele et al., 2022; Binkiewicz et al., 2017).
With the network structure established, we now relate the latent positions $X_{i\cdot}$ to the node-level observations $(T_{i},C_{i\cdot},Y_{i})$.
Assumption 3 (Linear Conditional Expectations)
The outcome regression functional is linear in $T_{i},C_{i\cdot}$, and $X_{i\cdot}$ and the mediator regression functional is linear in $T_{i}$ and $C_{i\cdot}$:
$\underbrace{\mathbb{E}\left[Y_{i}\,\middle|\,T_{i},C_{i\cdot},X_{i\cdot}\right]}_{\mathbb{R}}$ $=\underbrace{\beta_{0}}_{\mathbb{R}}+\underbrace{T_{i}}_{\{0,1\}}\underbrace{\beta_{\text{t}}}_{\mathbb{R}}+\underbrace{C_{i\cdot}}_{\mathbb{R}^{1\times p}}\underbrace{\beta_{\text{c}}}_{\mathbb{R}^{p}}+\underbrace{X_{i\cdot}}_{\mathbb{R}^{1\times d}}\underbrace{\beta_{\text{x}}}_{\mathbb{R}^{d}},$ (outcome model)
$\underbrace{\mathbb{E}\left[X_{i\cdot}\,\middle|\,T_{i},C_{i\cdot}\right]}_{\mathbb{R}^{1\times d}}$ $=\underbrace{\theta_{0}}_{\mathbb{R}^{1\times d}}+\underbrace{T_{i}}_{\{0,1\}}\underbrace{\theta_{\text{t}}}_{\mathbb{R}^{1\times d}}+\underbrace{C_{i\cdot}}_{\mathbb{R}^{1\times p}}\underbrace{\Theta_{\text{c}}}_{\mathbb{R}^{p\times d}}$ (mediator model)
The columns of $T,C$ and $X$ must be linearly independent for regression coefficients to be identifiable. The latent dimension $d$ and the number of nodal controls $p$ are constants that do not vary with sample size.
In this model, $\beta_{0}$ and $\theta_{0}$ play the role of intercept terms, while $\beta_{\text{t}}$ and $\beta_{\text{c}}$ encode the average associations between nodal covariates $T,C$ and nodal outcomes $Y$, conditional on the effect of the latent positions $X$. $\theta_{\text{t}}$ and $\Theta_{\text{c}}$ describe how latent positions in the network vary with nodal covariates. In the general sub-gamma edge model that we consider here, it is difficult to provide an interpretation of $\beta_{\text{x}}$ and the mediator coefficients because the latent positions $X$ can take on several roles depending on the precise parametric sub-model under consideration (e.g., under the SBM as compared to the more general RDPG). Nonetheless, $\beta_{\text{x}}$ represents the average association between latent network structure and nodal outcomes, conditional on nodal covariates. Since $X$ is only identified up to orthogonal rotation, $\beta_{\text{x}}$ and $\Theta$ are similarly only identified up to orthogonal rotation.
When we combine the statistical model entailed by Assumption 3 with the previous counterfactual assumptions entailed by Assumption 1, the regression coefficients $\beta$ and $\Theta$ have known causal interpretations.
Proposition 3 (VanderWeele and Vansteelandt 2014)
Under Assumptions 1 and 3
$\Psi_{\mathrm{nde}}\left(t,t^{*}\right)$ $=\left(t-t^{*}\right)\,\beta_{\text{t}},\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (1)
$\Psi_{\mathrm{nie}}\left(t,t^{*}\right)$ $=\left(t-t^{*}\right)\,\theta_{\text{t}}\,\beta_{\text{x}}$ (2)
Remark 4
An immediate question is whether Proposition 3 is useful, given that $\theta_{\text{t}}$ and $\beta_{\text{x}}$ are subject to an unknown orthogonal transformation. Luckily, non-identifiability of the regression coefficients does not impact identifiability of causal estimands. $\Psi_{\mathrm{nde}}$ depends only on $\beta_{\text{t}}$, which is fully identified. On the other hand, $\Psi_{\mathrm{nie}}$ is a function of $\theta_{\text{t}}$ and $\beta_{\text{x}}$, and the unknown orthogonal transformations for these estimates cancel (a result that follows immediately from Theorem 9 below). This is because $\Psi_{\mathrm{nie}}$ depends fundamentally on the projection of $Y$ onto the span of $\{T,C,X\}$ and the projection of $X$ onto the span of $\{T,C\}$, rather than the precise bases of those subspaces.
Remark 5
Proposition 3 identifies the natural direct and indirect effects based solely on outcome and mediator regression models, and does not require a propensity score model. It is possible to construct more complicated estimators that additionally leverage the propensity score, and these estimators can be more robust and efficient than the product-of-coefficients estimator considered here (Nguyen et al., 2021; Tchetgen Tchetgen and Shpitser, 2012). We are primarily interested in network regression in this work, and leave exploration of other estimators to future work.
2.4 Causal mechanisms in latent space
A first challenge when considering mediation in a latent space is to understand how latent mediation works in the abstract. To characterize one possible form of latent mediation, we develop some intuition in the context of degree-corrected mixed-membership stochastic blockmodels, as in Example 2. Recall that, under this model, the latent position $X_{i\cdot}$ encodes the group memberships of node $i$, and the “popularity” of node $i$ (i.e., the propensity of node $i$ to form connections with other nodes) [^1] More precisely, under the degree-corrected mixed-membership stochastic blockmodel, $\mathbb{P}\left(A_{ij}=1\,\middle|\,Z,B\right)=Z_{i\cdot}BZ_{j\cdot}^{T}$, where $Z\in\mathbb{R}_{+}^{n\times d}$ is a matrix of popularity-scaled group membership weights, and $B$ is a positive semi-definite mixing matrix. In particular, under this model, $X=ZB^{1/2}$. The $B^{1/2}$ factor in $X$ can be slightly counter-intuitive at first. For the sake of intuition, one can conceptualize interventions as applying to the block memberships $Z$, while holding the mixing matrix $B$ constant. In Proposition 19, we show that any intervention on $Z$ satisfying the parametric constraints of Assumption 3 is equivalent to an intervention on $X$, which also satisfies the parametric constraints of Assumption 3. Thus, for concreteness, one can think about interventions applying directly to the block memberships $Z$. Under Assumption 3, interventions can also simultaneously influence $B$, but these interventions may feel less intuitive.. The latent position $X_{i\cdot}$ is involved in several causal pathways.
  1. 1.
    The homophily-inducing pathway $T_{i}\to X_{i\cdot}$. Since $X_{i\cdot}$ encodes the group membership and popularity of node $i$, intervening on $X_{i\cdot}$ can cause node $i$ to participant in different communities, or change its popularity, or simultaneously translate node $i$ to new communities while also scaling its popularity. Assumption 3 implies that intervention must, on average, cause a translation in the latent space.
  2. 2.
    The network-formation pathway $X_{i\cdot}\to A_{i\cdot}$. Intervening on $X_{i\cdot}$ simultaneously modifies $\mathbb{P}\left(A_{ij}=1\,\middle|\,X\right)$ for all $j\in[n]$. That is, intervening on node $i$’s community membership changes node $i$’s probability of connecting to every other node.
  3. 3.
    The social outcome effect pathway $X_{i\cdot}\to Y_{i}$. This encodes the idea that community membership and popularity influence outcomes.
To develop intuition for interventions on the latent positions, consider a degree-corrected mixed membership stochastic blockmodel with five blocks, where nodes are primarily members of a single block and there is a small amount of degree heterogeneity. Interventions can increase or decrease partipication in particular communities. For instance, in Figure 3, we visualize an intervention that decreases participation in the second community while increasing participation in the third community, in the latent space.
Figure 3: Block memberships in a degree-corrected mixed-membership stochastic blockmodel before and after intervention. Each row corresponds to a node in the network and each column corresponds to a block in the blockmodel. Darker colors indicated increased participation in a given block. The intervention decreases participation in the second community slightly and increases participation in the third community more dramatically. Nodes are sorted by community membership.
Intervention in turn alters the probability of friendship between pairs of nodes, and thus leads to different counterfactual networks via the pathway $T\to X\to A$. To understand the impact of the intervention on the network itself, Figure 4 visualizes the difference between counterfactual networks where intervention does and does not occur. In the top left panel, we see friendship probabilities when $T$ does not impact $X$, and in the top right panel, we see friendship probabilities when $T$ causes a change in $X$. Intervention causes all nodes to participate in the third community. The mixing matrix $B$ is a diagonal matrix in this example, so nodes participating in the third community have some probability to connect with other nodes also participating in the third community. This means that, under intervention, all nodes in the network are more likely to connect with one another, due to mutual participation in the third block. The effect is especially pronounced for nodes that began in the third block. The bottom panels of Figure 4 show corresponding random samples from the networks conditional on latent positions $X$.
(a) $P$ without intervention.
(b) $P$ with intervention.
(c) Realized network without intervention
(d) Realized network with intervention
Figure 4: A visualization of how intervention impacts the probability of edge formation $P$, and ultimately the realized network $A$. In all cases matrices as visualized as heatmaps. Each element of the heatmap corresponds to one edge in the network. Sub-figure 4(a) shows edge formation probabilities pre-intervention, and Sub-figure 4(b) shows edge formation probabilities post-intervention. Sub-figure 4(c) shows a network realized when the intervention doesn’t occur, and Sub-figure 4(d) shows a network realized when the intervention does occur. Nodes are sorted by community membership.
This synthetic example is primarily to develop intuition about the latent positions, and downstream consequences of an intervention. Even with this intuition, it may be difficult to understand what it means to intervene on a latent position. In practice, we believe that the best way to resolve this ambiguity is by interpreting the latent positions in the context of the relevant data application. In practice, we compute estimates of the latent positions using the adjacency spectral embedding (see Section 3). With estimates of the latent positions in hand, it is often much easier to reason about causal mechanisms, because the estimated positions can be given a concrete interpretation via inspection, auxiliary data and domain knowledge. In the psychometric data application of Section 5.2, for instance, the latent positions have a very clear interpretation. To interpret the latent positions, we find that Gaussian mixture models and varimax rotation are often valuable tools (Rohe and Zeng, 2023; Rubin-Delanchy et al., 2022; Zhang et al., 2021; Priebe et al., 2019).
To summarize: we believe that the degree-corrected mixed-memberhip stochastic blockmodel offers compelling intuition about the structures that latent positions can represent. This intuition is helpful to determine when a latent mediation model may be applicable: for instance, in settings where one suspects that social groups act as mediators. However, there is an additional confirmatory step necessary for empirical work. It is critical to confirm that the estimated latent positions do in fact capture the hypothesized mediating constructs. Estimates of latent positions might reflect other network structure unrelated to hypothesized mediators. This is a weakness of our smoking example: we know from extensive sociology research that the adolescent social network is composed of numerous small social groups. However, we cannot confirm that the social structure observed by sociologists is the same as the social structure captured by the network principal components, an issue that we discuss in more detail in Section 5.1. In contrast, in the psychometric example of Section 5.2, the estimated latent positions are clearly measures of latent constructs of interest.
2.5 Relation to peer effects
The mediation mechanism that we have proposed considers how network effects manifest through group-level dynamics rather than individual-level interactions. In network settings, it is commonly assumed that there may be peer effects operating at the level of individual interactions. Two especially pertinent peer effects are contagion and interference (Hu et al., 2022; Ogburn et al., 2020). Contagion occurs when the outcome of neighbor $j$ impacts the outcome of node $i$ (i.e., there is a path $Y_{j}\to Y_{i}$). Interference occurs when the treatment applied to neighbor $j$ impacts the outcome of node $i$ (i.e., there is a path $T_{j}\to Y_{i}$). In the context of our smoking example, eliding some simultaneity issues, contagion would occur if a student’s tobacco usage were influenced by their friends’ tobacco usage. Interference would occur if a student’s tobacco usage were influenced by their friends’ sex.
Both our counterfactual assumptions and the parametric form of our outcome model assume that there are no peer effects. This restriction clearly limits the applicability of our method, but there are some caveats. First, in ongoing work, we are extending the network mediation model to allow for these peer effects. The extension is quite involved, but preliminary results and related work suggest that, at least some of the time, it is possible to estimate both mediated effects and peer effects using ordinary least squares and estimated latent positions $\widehat{X}$, as we propose here (Chang and Paul, 2024; Lee, 2002; Paul et al., 2022b; Trane, 2023). That is, one can view the network mediation model, and our characterization of mediation in a latent space, as descriptions of an important subset of causal mechanisms in networks. In many cases, these mechanisms should be jointly modelled, but developing methods to do so is a complicated technical endeavor that is outside the scope of this paper.
Part of the challenge in jointly modelling latent mediation and peer effects is that the mechanisms can be difficult or impossible to distinguish. Shalizi and Thomas (2011) showed this in a non-parametric longitudinal setting, but a very similar consideration arises in the parametric cross-sectional setting. To see this, consider the “linear-in-sums” model, one popular way to model contagion and interference (Paul et al., 2022a; McFowland and Shalizi, 2021; Egami and Tchetgen Tchetgen, 2021; Hu et al., 2022; Bramoullé et al., 2020):
$\mathbb{E}\left[Y_{i}\,\middle|\,T_{i},C_{i\cdot},X,A,Y_{-i}\right]=\beta_{0}+T_{i}\beta_{\text{t}}+C_{i\cdot}\beta_{\text{c}}+X_{i\cdot}\beta_{\text{x}}+\beta_{\text{Ay}}\sum_{j\neq i}^{n}A_{ij}Y_{j}+\beta_{\text{At}}\sum_{j\neq i}^{n}A_{ij}T_{j}.$
The coefficient $\beta_{\text{Ay}}$ encodes a contagion effect, and the coefficient $\beta_{\text{At}}$ encodes an interference effect. In forthcoming work, we show that columns of the design matrix corresponding to $\beta_{\text{x}}$, $\beta_{\text{Ay}}$ and $\beta_{At}$ are collinear in the asymptotic limit. That is, the interference column $AT$ and the contagion column $AY$ of the design matrix are both contained in the column space of $X$ in large samples. This occurs because the column spaces of $A$ and $X$ get closer with increasing sample size under the sub-gamma model, and thus peer effects that are linear functions of $A$ can be equivalently represented as linear functions of $X$. From a parametric perspective, the coefficient $\beta_{\text{x}}$ generalizes “linear-in-sums” style peer effects. In the context of our smoking example, the causal effect of belonging to a particular friend group is asymptotically indistinguishable from a contagion effect that depends on the number of friends who are smokers. Thus, the causal effects along the pathway $X_{i\cdot}\to Y_{i}$ can be thought of as effects of group memberships, or alternatively, as consequences of diffusions over the network. The fact that this equivalence holds only in the asymptotic limit, however, introduces a number of subtle caveats, which we describe in a forthcoming manuscript.
In addition to “linear-in-sums” style peer effect, it is also possible to consider “linear-in-means” style peer effects,
$\mathbb{E}\left[Y_{i}\,\middle|\,T_{i},C_{i\cdot},X,A,Y_{-i}\right]=\beta_{0}+T_{i}\beta_{\text{t}}+C_{i\cdot}\beta_{\text{c}}+X_{i\cdot}\beta_{\text{x}}+\beta_{\text{Ay}}\sum_{j\neq i}^{n}\frac{A_{ij}}{d_{i}}Y_{j}+\beta_{\text{At}}\sum_{j\neq i}^{n}\frac{A_{ij}}{d_{i}}T_{j}$
where $d_{i}=\sum_{j}A_{ij}$ is the degree of node $i$ (Bramoullé et al., 2009). In the linear-in-means approach, $\beta_{\text{x}}$, $\beta_{\text{Ay}}$ and $\beta_{At}$ are identified, even in the asymptotic limit, under some non-trivial assumptions about the network structure. Recent work has considered these models from a statistical point of view (Chang and Paul, 2024; Lee, 2002; Paul et al., 2022b; Trane, 2023), although we believe there are still substantial identifications concerns to address, from both a causal and a statistical perspective.
3 Estimation theory for principal components network regression
Having established a model, we have several estimation targets. First, there are the network regression coefficients $\beta$ and $\Theta$ from Assumption 3. Once we estimate the coefficients $\beta$ and $\Theta$, we can plug them into Equations (1) and (2) to obtain estimates of the natural direct and indirect effects, respectively.
3.1 Principal components network regression
Before discussing estimation, we simplify notation by collecting the node-level covariates into $W=\begin{bmatrix}1&T&C\end{bmatrix}\in\mathbb{R}^{n\times(p+2)}$. Assumption 3 can then be re-written as
$\mathbb{E}\left[Y_{i}\,\middle|\,W_{i\cdot},X_{i\cdot}\right]$ $=W_{i\cdot}\beta_{\text{w}}+X_{i\cdot}\beta_{\text{x}},$ (outcome model) (3)
$\mathbb{E}\left[X_{i\cdot}\,\middle|\,W_{i\cdot}\right]$ $=W_{i\cdot}\Theta.$ (mediator model)
We would like to estimate $\beta_{\text{w}},\beta_{\text{x}}$ and $\Theta$ by applying ordinary least squares regression to the vertex-level latent positions $X$. Unfortunately, the latent positions $X$ are unobserved. To contend with this, we will estimate $X$ from the observed network, then plug in $\widehat{X}$ for $X$ in subsequent regressions. We use the adjacency spectral embedding (ASE; Sussman et al., 2012) to achieve this estimation, but we note that several other methods are available (Xie and Xu, 2020, 2021; Wu and Xie, 2022).
Definition 6 (ASE; Sussman et al. (2014))
Given a network with adjacency matrix $A$, the $d$-dimensional adjacency spectral embedding (ASE) of $A$ is defined as
$\widehat{X}=\widehat{U}\widehat{S}^{1/2}\in\mathbb{R}^{n\times d},$
where $\widehat{U}\widehat{S}\widehat{V}^{T}$ is the rank-$d$ truncated singular value decomposition of $A$. That is, $\widehat{S}\in\mathbb{R}^{d\times d}$ is diagonal, with entries given by the $d$ leading singular values of $A$, and $\widehat{U},\widehat{V}\in\mathbb{R}^{n\times d}$ have the corresponding $d$ orthonormal singular vectors as their columns.
The spectral embeddings $\widehat{X}$ converge to the true $X$ uniformly over the rows of $X$, up to orthogonal non-identifiability (Levin et al., 2022; Lyzinski et al., 2014), suggesting that ordinary least squares estimates based on $\widehat{X}$ rather than $X$ may have nice asymptotic behavior. We thus construct least squares estimators for the regression coefficients as follows.
Definition 7 (Regression Point Estimators)
Define $\widehat{D}=\begin{bmatrix}W&\widehat{X}\end{bmatrix}\in\mathbb{R}^{n\times(2+p+d)}$. We estimate $\beta_{\text{w}}$ and $\beta_{\text{x}}$ via ordinary least squares as follows
$\begin{bmatrix}\widehat{\beta}_{\text{w}}\\ \widehat{\beta}_{\text{x}}\end{bmatrix}=\left(\widehat{D}^{T}\widehat{D}\right)^{-1}\widehat{D}^{T}Y.$
Similarly, we estimate $\Theta$ via ordinary least squares as
$\widehat{\Theta}=\left(W^{T}W\right)^{-1}W^{T}\widehat{X}.$
Definition 8 (Regression variance estimators)
With notation as above, we define covariance estimators
$\widehat{\Sigma}_{\beta}$ $=\widehat{A}_{\beta}^{-1}\cdot\widehat{B}_{\beta}\cdot\left(\widehat{A}_{\beta}^{-1}\right)^{T}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }$
$\widehat{\Sigma}_{\operatorname*{vec}\left(\Theta\right)}$ $=\widehat{A}_{\operatorname*{vec}\left(\Theta\right)}^{-1}\cdot\widehat{B}_{\operatorname*{vec}\left(\Theta\right)}\cdot\left(\widehat{A}_{\operatorname*{vec}\left(\Theta\right)}^{-1}\right)^{T},$
where $I_{d}$ is a $d\times d$ identity matrix, and letting $\widehat{\xi}_{i\cdot}=\widehat{X}_{i\cdot}-W_{i\cdot}\widehat{\Theta}$, we define
$\widehat{A}_{\beta}=\frac{\widehat{D}^{T}\widehat{D}}{n},$ $\widehat{B}_{\beta}=\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}-\widehat{D}_{i\cdot}\widehat{\beta}\right)^{2}\widehat{D}_{i\cdot}^{T}\widehat{D}_{i\cdot},$
$\widehat{A}_{\operatorname*{vec}\left(\Theta\right)}=\frac{I_{d}\otimes W^{T}W}{n},\quad\text{and}$ $\widehat{B}_{\operatorname*{vec}\left(\Theta\right)}=\frac{1}{n}\sum_{i=1}^{n}\widehat{\xi}_{i\cdot}^{T}\,\widehat{\xi}_{i\cdot}\otimes W_{i\cdot}^{T}W_{i\cdot}.$
Our main technical results state that the ordinary least squares estimates based on $\widehat{X}$ converge to the same asymptotic distribution as the ordinary least squares estimates based on $X$. That is, the estimated regression coefficients asymptotically behavior as if we had access to the true latent positions, even though we do not. In order for plug-in estimation to be useful, the estimates based on the true latent positions $X$ must themselves be well-behaved. Under fairly weak conditions, estimates based on the true latent positions are asymptotically normal.
Assumption 4 (Regularity Conditions for Ordinary Least Squares $M$-estimation)
We require standard regularity conditions for ordinary least squares estimates for both the outcome model and the mediator model. See Chapter 7 of Boos and Stefanski (2013) and Chapter 5 of van der Vaart (1998) for additional discussion.
  1. 1.
    Define $\xi=X-\mathbb{E}\left[X\,\middle|\,W\right]\in\mathbb{R}^{n\times d}$ to be the matrix of errors in the mediator regression. $\xi_{ij}$ and $\xi_{i^{\prime}j}$ are independent for $i\neq i^{\prime}$ and all $j\in[d]$, and that $\mathbb{E}\left[\xi_{ij}\,\middle|\,W\right]=0$ for all $i\in[n],j\in[d]$. Further,
    $A_{\operatorname*{vec}\left(\Theta\right)}\equiv\mathbb{E}\left[\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}W_{i\cdot}^{T}W_{i\cdot}\right]\in\mathbb{R}^{p\times p}$
    exists and is non-singular, and the matrix $B_{\operatorname*{vec}\left(\Theta\right)}\in\mathbb{R}^{d\times d}$ defined according to
    $[B_{\operatorname*{vec}\left(\Theta\right)}]_{j,j^{\prime}}\equiv\mathbb{E}\left[\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\xi_{ij}\,W_{i\cdot}^{T}W_{i\cdot}\,\xi_{ij^{\prime}}\right]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ j,j^{\prime}\in[d]$
    exists and has all entries finite.
  2. 2.
    Define $\widehat{D}=\begin{bmatrix}W&\widehat{X}\end{bmatrix}\in\mathbb{R}^{n\times(2+p+d)}$, let $\varepsilon=Y-\mathbb{E}\left[Y\,\middle|\,D\right]\in\mathbb{R}^{n}$ be the vector of errors in the outcome regression. The $\varepsilon_{i}$ are independent, and obey $\mathbb{E}\left[\varepsilon_{i}\,\middle|\,D\right]=0$ for all $i\in[n]$. Further, $A_{\beta}\equiv\mathbb{E}\left[\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}D_{i\cdot}^{T}D_{i\cdot}\right]\in\mathbb{R}^{(p+d)\times(p+d)}$ exists and is non-singular, and
    $B_{\beta}\equiv\mathbb{E}\left[\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\,D_{i\cdot}^{T}D_{i\cdot}\right]$
    exists and is finite.
For consistent estimation of $\beta$ and $\Theta$ based on known $X$, the key requirement is that the residuals $\varepsilon_{i}$ and $\xi_{ij}$ have (conditional) mean zero, given the covariates $D$ and $W$, respectively. Fundamentally, this means that $\mathbb{E}\left[Y_{i}\,\middle|\,W_{i\cdot},X_{i\cdot}\right]$ must be linear in both $W_{i\cdot}$ and $X_{i\cdot}$ and that $\mathbb{E}\left[X_{i\cdot}\,\middle|\,W_{i\cdot}\right]$ must be linear in $W_{i\cdot}$. We note that the errors $\varepsilon$ and $\xi$ need not come from a particular distribution: any distribution satisfying the assumptions on $A_{\operatorname*{vec}\left(\Theta\right)},B_{\operatorname*{vec}\left(\Theta\right)},A_{\beta}$, and $B_{\beta}$ will do. The rows of $\xi$ must be independent of one another, but arbitrary dependence is acceptable within the rows of $\xi$. Once the regularity conditions in Assumption 4 are satisfied, these least-squares estimates for $\beta$ and $\Theta$ are asymptotically normal about their estimands. Further, the well-known “robust” or “sandwich” covariance estimator is a consistent estimator for the covariance structures of these asymptotic distributions and the ordinary least squares estimates and covariance estimators can be used to obtain asymptotically valid confidence intervals for $\beta$ and $\Theta$.
Since we do not have access to the latent positions $X$, we must estimate them using the adjacency matrix, and as a result we need additional structure beyond Assumption 4.
Assumption 5 (Conditions for Two-stage Estimability)
With notation as above, we assume that
  1. 1.
    The smallest non-zero eigenvalue $\lambda_{d}$ of $P$ grows at a sufficiently fast rate as a function of $n$ relative to the noise parameters $\nu_{n}$ and $b_{n}$. In particular,
    $\left(\nu_{n}+b_{n}^{2}\right)=o\left(\frac{\lambda_{d}^{2}}{n\log^{2}n}\right).$
  2. 2.
    The entries of the observation error vector $\varepsilon$ have bounded second moments, i.e., $\max_{i\in[n]}\mathbb{E}\left[\varepsilon_{i}^{2}\right]<B$ for $B>0$ not depending on $n$. Further, we assume that the error vector is such that $n^{-1}\sum_{i}\mathbb{E}\left[\varepsilon_{i}^{4}\right]$ is bounded as a function of $n$.
  3. 3.
    The rows $W_{1\cdot},W_{2\cdot},\dots,W_{n\cdot}$ of $W$ are independent. Within each row $W_{i,\cdot}$, the elements may be dependent, but are marginally sub-Gaussian with fixed, shared parameter $\sigma>0$.
  4. 4.
    The latent positions $X_{1},X_{2},\dots,X_{n}$ are such that $n^{-1}\sum_{i}\mathbb{E}\left[\left\lVert\xi_{i\cdot}\right\rVert^{2}\right]$ and $n^{-1}\sum_{i}\mathbb{E}\left[\left\lVert X_{i\cdot}\right\rVert^{4}\right]$ are bounded as functions of $n$.
The first condition is a sufficient condition for $\widehat{X}$ to concentrate around $X$ (see Levin et al., 2022, for further discussion). This condition primarily places requirements on the density of edges in the network, and in particular requires that the expected average degree is $\omega(\sqrt{n}\log n)$. In a random dot product graph, all degrees are asymptotically of the same order as the average degree, so our results are for dense networks. See Remark 12 for some additional comments on sparsity.
The second condition puts some additional (weak) conditions on the error in the outcome regression, beyond those already required for $M$-estimation. While the second condition strengthens Assumption 4, it still makes no distributional assumptions. For example, if the $\varepsilon_{i}$ are independent and identically distributed sub-gamma random variables, our bounded second moment condition is satisfied. The third condition is necessary for control over $\left\lVert W\right\rVert$ in our proofs. The fourth-moment assumptions in the second and fourth conditions are needed to ensure convergence of our covariance estimator based on the spectral estimates. They ensure that error terms between this spectral-based covariance estimate and the covariance estimate based on the true (but unobserved) latent positions are suitably close. We anticipate that the moment bounds in the second, third and fourth conditions can be relaxed, but at the expense of additional proof complexity.
Our theoretical results, in their most general form, require certain growth rates on the sub-gamma parameters $\nu_{n}$ and $b_{n}$, and the largest and smallest non-zero eigenvalues $\lambda_{1}$ and $\lambda_{d}$ of $P$, described in Assumption 6. For ease of presentation, we note that these rates are satisfied by random dot product graphs (and therefore stochastic blockmodels, since stochastic blockmodels are a submodel of random dot product graphs) and present Theorem 9 in the setting of a random dot product graph. Fully general bounds in terms of $n,\nu_{n},b_{n},\lambda_{1}$ and $\lambda_{d}$ along with proof details may be found in the Appendix.
Theorem 9
If Example 1, and Assumptions 34, and 5 hold, then there exists a sequence of orthogonal matrices $\{Q_{n}\}_{n=1}^{\infty}$ such that
$\sqrt{n}\,\widehat{\Sigma}_{\operatorname*{vec}\left(\Theta\right)}^{-1/2}\begin{pmatrix}\operatorname*{vec}\left(\widehat{\Theta}\,Q_{n}^{T}\right)-\operatorname*{vec}\left(\Theta\right)\end{pmatrix}$ $\to\mathcal{N}\left(0,I_{pd}\right),\text{ and }$
nΣ^β1/2(β^wβw